40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  OMC elog  Not logged in ELOG logo
Message ID: 115     Entry time: Wed Apr 17 07:30:04 2013
Author: Koji 
Type: Optics 
Category: General 
Subject: QPD path glued 

Yesterday, all of the glass components for the QPD path were glued.

- Check the alignment of the beam with the cavity.

- Placed the prisms

- Placed the QPD mount for the gluing test. An alignment disk instead of a diode was placed on the mount.

- Checked the spot positions at the QPDs. A CCD camera with a lens was used to find the spot.
  The spots were ~0.5mm lower on the QPD1, and ~1mm lower on the QPD2.

- Glued the first steering mirror while the spot position was continuously monitored.

- Glued the BS in the QPD path while the spot position was monitored.

- FInally a glass bracket was glued.

- The spot on QPD2 was too low to absorb by the QPD shim.

- Once the steering mirror was clamped by a cantilever spring (to prevent slipping), the spot moved up a bit.
  (Or, we should say, the cantilever misaligned the optics a bit in pitch in a preferrable direction.)

- The other steering mirror is clamped by a cantilever spring (to prevent slipping), the spot moved up a bit.
  Or, we should say, the cantilever misaligned the optics a bit in pitch in a preferrable direction.)

- The last steering mirrors was also glued in a same way. As a result the spot is 0.5mm below the center of the alignment disk.

- Once the PD mounting brackets were glued, we can't place the QPD mount on it as the PEEK bar can't be inserted without moving the gluing template.

- The QPD mount with out the glass bracket was used to check the alignment of the beam dumps.
  As the beam dumps have a wide aperture and the yaw alignment of the QPD is big, we could accommodate the beams in the dumps easily.

- The dumps were glued.

ELOG V3.1.3-